# Parameter control for eccentric, precessing binary black hole simulations with SpEC

### v1

@@ II.3 Eccentric Post-Newtonian Dynamics
We fit the numerical data with the Keplerian equations of eccentric motion augmented with post-Newtonian corrections as given in Ref. Poisson and Will . We use these post-Newtonian equations to extract the Keplerian parameters $\vec{\theta}_{\rm orb}$ from numerical data and to compute the inverse operator ${\cal{G}}$ in Eq. (3).
A generic Keplerian orbit at Newtonian order, 0PN, is characterized by its semi-major axis $a$, eccentricity $e$, and mean anomaly $\ell$ defined at some reference time. These three quantities are related via Kepler’s equations: $\bar{\Omega}t=u(t)-e\sin{u(t)}$ and $\bar{\Omega}=\sqrt{M/a^{3}}$. The eccentric anomaly, $u(t)$, is obtained by inverting Kepler’s equation, approximated as: $u(t)=\bar{\Omega}t+\ell$. To next, 1PN order, the Keplerian parameters and all equations obtain corrections. These can be expressed in closed form and as a function of the Newtonian orbital parameters $(a,e,\ell)$. We present these equations and their derivation in Appendix A. The pertinent point is that there is a closed-form expressions for the (derivative of the) orbital angular frequency as a function of time, $\dot{\Omega}(t;a,e,\ell).$
@@ II.4 PN Trajectory Fitting Model
In the post-Newtonian framework, $\dot{\Omega}(t)$ deviates from zero (circular orbit) due to three effects:
@@
The second term in Eq. (5) encodes 2PN spin-spin interactions Buonanno et al. (2011):
@@
The final term in Eq. (5) encodes eccentric effects, therefore it is a function of the orbital parameters. To 1PN order Poisson and Will
@@
Appendix A.2 also presents a derivation of Eq. (8) based on Ref. Poisson and Will . Crucially, $\delta\dot{\Omega}_{e}(t)$ at this order depends only on $(a,e,\ell)$, so we can use this expression to fit numerical data and obtain estimates for the orbital parameters.
Motivated by the three terms of Eq. (5), we define the following model to fit the numerical data, shown by the dashed curve in Fig. 1:
@@
The first line of Eq. (9) captures radiation-reaction. It depends on amplitude parameters $C_{1}$ and $C_{2}$ that we fit for. The coalescence time, $T_{c}$, is calculated from the evolution data using using a 0PN approximation for the time to merger:
@@ II.5 Fitting the Numerical Data
Next, we must decide how long to evolve the BBH after $t_{T}$. We motivate this choice with three arguments. First, the fitting window must be short to mitigate computational cost for the trial simulation whose main goal is to evaluate Eq. (3) and provide improved initial parameters. Second, the eccentricity, semi-major axis, and mean anomaly are PN parameters that evolve due to radiation reaction and become ill-defined close to merger. The fitting window must therefore both be short and as far away from merger as possible. Third, since we are fitting harmonic functions, at least a few orbits are required to achieve enough precision. In practice, we evolve the trial simulation for a little less than 5 orbits up to a final time $T_{\rm end}=10\pi/\Omega_{0}$, calculated based on the initial orbital angular frequency. The PN fit then occurs between times $t_{T}$ and $T_{\rm end}$.
@@ II.6.2 Spin
Figure 4: Application of the parameter control loop outlined in Fig. 2 for an equal-mass binary with vanishing spins and target orbital parameters $\vec{\theta}_{\rm orb}{{}_{,T}}=\{a_{T}=15\,M,e_{T}=0.2,\ell_{T}=\pi\}$. The top left panel shows the NR data for each iteration of the control loop (top subpanel) and the residuals between the PN fit and NR data for each iteration (bottom subpanel). The rightmost grey vertical line indicates $t=t_{T}$. The PN fitting window occurs between the two vertical gray lines. The difference between the achieved, i.e., fitted from the NR data, orbital parameters in each iteration $i$ and the target parameters is plotted in the remaining panels: eccentricity (top right), semi-major axis (bottom left), and mean anomaly (bottom right). All parameters converge to their target value after $3$ iterations
@@
III Eccentric, non-precessing orbits
We begin with an example of the parameter control process. Figure 4 shows results for an equal-mass binary, target orbital parameters $\vec{\theta}_{\rm orb}{{}_{,T}}=\{a_{T}=15\,M,e_{T}=0.2,\ell_{T}=\pi\}$, and vanishing spins. [^1] Unless otherwise indicated, we set $\ell_{T}$ to apastron as this choice leads to more stable evolutions. The top left panel shows the numerical data for $\dot{\Omega}(t)$ for each of the three iterations required to achieve the eccentricity threshold. Each iteration follows the control process shown in Fig. 2 to produce fitted orbital parameters for the numerical data and update the initial data parameters of the next iteration accordingly.
Figure 5: Impact of target eccentricity value on parameter convergence. We plot the difference between the fitted and target eccentricity (left panel) and semi-major axis (right panel) at each iteration for binaries with equal masses, vanishing spins, $a_{T}=15\,M$ (top) or $r_{a,T}=60\,M$ (bottom), and different values of target eccentricity $e_{T}$ (color bar). We achieve parameter convergence for all the cases we attempted, including eccentricities up to $0.65$. For $e_{T}>0.65$, the BHs risk head-on collision, which terminates the control loop prematurely before obtaining $\vec{\theta}_{ID,T}$. Thus we omit cases $e_{T}>0.65$ in this work. As $e_{T}$ increases to 0.65, more iterations are required for parameter convergence. We also inlay separation over time plots based on the final (most converged) iteration of the control loop.
The remaining panels show the difference between the achieved (fitted from the NR data) and the target eccentricity (top right), semi-major axis (bottom left), and mean anomaly (bottom right) as a function of iteration. Here and in subsequent figures, $a^{(i)},e^{(i)}$ and $\ell^{(i)}$ refer to the fitted orbital parameters in each iteration $i$ of the control loop. Even though we only impose a tolerance threshold on the eccentricity, all orbital parameters converge to their target values after $3$ iterations. In fact, the mean anomaly, $\ell$, converges almost immediately. Since this behavior typical in subsequent explorations, we omit the mean anomaly from now on. Moreover, the semi-major axis converges in 1 fewer iteration than the eccentricity. This justifies continuing to consider only the eccentricity in the parameter control tolerance.
@@ III.1 Impact of target eccentricity value
Figure 5 shows parameter convergence at fixed $a_{T}$ and $e_{T}<0.5$ (top) and fixed $r_{a,T}$ and $e_{T}\geq 0.5$ (bottom). We show the difference between the target and fitted eccentricity (left panels) and semi-major axis (right panels) as a function of iteration and colored by the target eccentricity value. The inlaid plots show the BH coordinate separation over time, using the same color scheme. In the top row, the separation reflects a constant $a_{T}$ for each $e_{T}<0.5$. As the eccentricity increases (darker purple), the average separation stays constant but the apastron and periastron passages become more extreme. In the bottom row, all trajectories begin at the same apastron and result in roughly similar periastron values.
In all cases, i.e., for target eccentricty up to $e_{T}=0.65$, the eccentricity and semi-major axis converge to their target value, with higher eccentricities requiring on average more iterations. This is likely because higher eccentricity orbits are more relativistic, thus the post-Newtonian approximation to the dynamics is less accurate. Such less accurate fitting equations lead to less efficient initial data parameter updates in each control loop iteration, thus requiring more iterations to converge.
@@ III.2 Aligned spins
Figure 6: Impact of aligned spins on parameter control. We plot the difference between the fitted and target eccentricity (left panel) and semi-major axis (right panel) at each iteration for binaries with equal masses, $a_{T}=15\,M$, $e_{T}=\{0.1,0.2,0.3\}$ (top to bottom) and different values of the aligned spin (color bar). We successfully converge to the target parameters for eccentric, aligned spin binaries, with higher eccentricities again requiring more iterations.
Figure 7: Comparing the simultaneous (top) and the sequential (bottom) parameter control approaches. We consider equal-mass binaries with $a_{T}=15\,M$, $e_{T}=0.1$, and different values of $\chi_{1,x,T}$ (colorbar). We show the difference between the target and the achieved value for the eccentricity (left), semi-major axis (middle), and in-plane spin angle, $\arccos(\vec{\chi}_{T}\cdot\vec{\chi}^{(i)})$ (right). In the sequential case, we plot results from the second parameter control loop, i.e., after obtaining a simulation with the target orbital parameters and no in-plane spins. Both methods achieve convergence, though as expected the simultaneous method needs more iterations. Moreover, the orbital parameters are closer to their target value in the first iteration for the sequential method (bottom) and the spin direction converges monotonically (right).
Next, we explore the impact of aligned spins. Aligned spins remain constant in both direction and magnitude (modulo horizon absorption effects Scheel et al. (2015)). Therefore the rotations of Fig. 3 reduce to the identity, and the spins are not updated during the eccentricity-control iterations. However, the spins still affect the binary dynamics and rate of orbital decay and could therefore impact the orbital parameter convergence. We study binaries with equal masses, $a_{T}=15\,M$, $e_{T}=\{0.1,0.2,0.3\}$, and vary the amount of aligned spin $\chi_{z,A,T}=\chi_{z,B,T}$ in $[-0.7,0.7]$. Results are shown in Fig. 6 for the different target eccentricities (top to bottom) and aligned spin values (colorbar). In all cases, the eccentricity and semi-major axis converge to their target value after $\leq 7$ iterations.
@@ IV Eccentric, precessing orbits
In this section, we generalize to precessing binaries. Going forward, we assume one spinning BH and no aligned component for simplicity. In-plane spins precess in the inertial frame, causing the BH frame to change orientation. The parameter control loop of Fig. 2 and the updating formulas in Eqs. (11)-(13) and (14) treat the orbital parameters and spins independently for convenience. Since the orbital parameters and the spins are updated independently, we begin by exploring two parameter control methods:
@@
    Simultaneous. We follow the process of Fig. 2 where in the first iteration we simulate an eccentric, precessing binary and then iteratively update the orbital parameters and the spins, per Eqs. (11)-(13) and (14).
@@
Figure 7 compares parameter convergence for the simultaneous (top) and the sequential (bottom) methods for an equal-mass binary, $a_{T}=15\,M$, $e_{T}=0.1$ and different values of $\chi_{1,x,T}$, and with $\chi_{1,y,T}=\chi_{1,z,T}=|\vec{\chi}_{2,T}|=0$. For the sequential case, we present only the second parameter control loop, i.e., the one that updates both the orbital parameters and the spins. Beyond the eccentricity (left) and the semi-major axis (middle), we show the angle between the target and the actual spin at $t_{T}$ (right). As before, all parameter control loops use a tolerance threshold based solely on eccentricity.
@@
Figure 8: Parameter convergence for precessing spin and varying spin magnitudes. We plot the difference between the target and the fitted eccentricity (left), semi-major axis (middle), and spin direction (right) for different values of the magnitude and direction of the in-plane spin. We choose $|\vec{\chi}_{1,T}|=0.3$ (solid) and $|\vec{\chi}_{1,T}|=0.6$ (dashed). The angle between the target in-plane spin and the $x$-axis is denoted as $\arccos(\vec{\chi}_{x}\cdot\vec{\chi}_{T})$ (colorbar). In all cases, the parameters converge after $\leq 6$ iterations with behavior depending more on the in-plane spin magnitude than the in-plane spin angle, c.f. also the top panel of Fig. 7, which is similar to this plot except it compares different in-plane spin magnitudes with the same in-plane spin angle.
Comparing the rate of parameter convergence between the two methods, the sequential method results in orbital parameters closer to their target values at the first iteration. This is evident for both the eccentricity and semi-major axis. Even though the second control loop in the sequential method plotted here starts from pre-converged orbital parameters, they still differ from their target value. This is because the first loop achieved the target parameters under zero in-plane spins, which are added in the second loop. Therefore the difference between the actual and the target eccentricity and semimajor axis in the first iteration is a consequence of the impact of in-plane spins on orbital dynamics.
Simultaneously adjusting the orbital parameters and the spin directions (top) results in spins that are no longer monotonically converging to their target value. Between the first and the second iteration, and even though both the eccentricity and the semi-major axis are closer to their target values, the spin angle difference increases from $\sim 5-10\,$deg to $\sim 20-30\,$deg. This is again a consequence of the coupling of orbital and spin dynamics. The spin is updated per Eq. (14) for the second iteration assuming the orbit of the first iteration. Despite this, in the third iteration where the orbit is not substantially updated, the spin angle difference quickly decreases. Overall, we achieve differences between the target and the actual spin angle ${\cal{O}}(\rm{deg})$, even though spins are not considered in the tolerance threshold.

### v2

@@ II.3 Eccentric post-Newtonian dynamics
We fit the numerical data with the Keplerian equations of eccentric motion augmented with post-Newtonian corrections as given in Ref. Poisson and Will (2014). We use these post-Newtonian equations to extract the Keplerian parameters $\vec{\theta}_{\rm orb}$ from numerical data and to compute the inverse operator ${\cal{G}}$ in Eq. (3).
A generic Keplerian orbit at Newtonian order, 0PN, is characterized by its semimajor axis $a$, eccentricity $e$, and mean anomaly $\ell$ defined at the epoch $t=0$. These three quantities are related via Kepler’s equations: $\bar{\Omega}t=u(t)-e\sin{u(t)}-\ell$ and $\bar{\Omega}=\sqrt{M/a^{3}}$. The eccentric anomaly, $u(t)$, is obtained by inverting Kepler’s equation, which in the small-eccentricity limit can be approximated as: $u(t)=\bar{\Omega}t+\ell$. To next, 1PN order, the Keplerian parameters and all equations obtain corrections. These can be expressed in closed form and as a function of the Newtonian orbital parameters $(a,e,\ell)$. We present these equations and their derivation in Appendix A. The pertinent point is that there is a closed-form expressions for the (derivative of the) orbital angular frequency as a function of time, $\dot{\Omega}(t;a,e,\ell).$
@@ II.4 PN trajectory fitting model
In the post-Newtonian framework, $\dot{\Omega}(t)$ deviates from zero (circular orbit) due to three effects
@@
The second term in Eq. (5) encodes 2PN spin-spin interactions Buonanno et al. (2011)
@@
The final term in Eq. (5) encodes eccentric effects, therefore it is a function of the orbital parameters. To 1PN order Poisson and Will (2014)
@@
Appendix A.2 also presents a derivation of Eq. (8) based on Ref. Poisson and Will (2014). Crucially, $\delta\dot{\Omega}_{e}(t)$ at this order depends only on $(a,e,\ell)$, so we can use this expression to fit numerical data and obtain estimates for the orbital parameters.
Motivated by the three terms of Eq. (5), we define the following model to fit the numerical data, shown by the dashed curve in Fig. 1
@@
The first line of Eq. (9) captures radiation-reaction. It depends on amplitude parameters $C_{1}$ and $C_{2}$ that we fit for. The coalescence time, $T_{c}$, is calculated from the evolution data using using a 0PN approximation for the time to merger
@@ II.5 Fitting the numerical data
Next, we must decide how long to evolve the BBH after $t_{T}$. We motivate this choice with three arguments. First, the fitting window must be short to mitigate computational cost for the trial simulation whose main goal is to evaluate Eq. (3) and provide improved initial parameters.
Second, the eccentricity, semimajor axis, and mean anomaly are PN parameters that evolve due to radiation reaction and become ill defined close to merger. The fitting window must therefore both be short and as far away from merger as possible. Third, since we are fitting harmonic functions, at least a few orbits are required to achieve enough precision. In practice, we evolve the trial simulation for a little less than 5 orbits up to a final time $T_{\rm end}=10\pi/\Omega_{0}$, calculated based on the initial orbital angular frequency. The PN fit then occurs between times $t_{T}$ and $T_{\rm end}$.
@@ II.6.2 Spin
Figure 4: Application of the parameter control loop outlined in Fig. 2 for an equal-mass binary with vanishing spins and target orbital parameters $\vec{\theta}_{\rm orb}{{}_{,T}}=\{a_{T}=15\,{\rm M},e_{T}=0.2,\ell_{T}=\pi\}$. The top left panel shows the NR data for each iteration of the control loop (top subpanel) and the residuals between the PN fit and NR data for each iteration (bottom subpanel). The rightmost gray vertical line indicates $t=t_{T}$. The PN fitting window occurs between the two vertical gray lines. The difference between the achieved, i.e., fitted from the NR data, orbital parameters in each iteration $i$ and the target parameters is plotted in the remaining panels: eccentricity (top right), semimajor axis (bottom left), and mean anomaly (bottom right). All parameters converge to their target value after $3$ iterations
@@
III Eccentric, nonprecessing orbits
We begin with an example of the parameter control process. Figure 4 shows results for an equal-mass binary, target orbital parameters $\vec{\theta}_{\rm orb}{{}_{,T}}=\{a_{T}=15\,{\rm M},e_{T}=0.2,\ell_{T}=\pi\}$, and vanishing spins. [^1] Unless otherwise indicated, we set $\ell_{T}$ to apastron as this choice leads to more stable evolutions. The top left panel shows the numerical data for $\dot{\Omega}(t)$ for each of the three iterations required to achieve the eccentricity threshold. Each iteration follows the control process shown in Fig. 2 to produce fitted orbital parameters for the numerical data and update the initial data parameters of the next iteration accordingly.
Figure 5: Impact of target eccentricity value on parameter convergence. We plot the difference between the fitted and target eccentricity (left panel) and semimajor axis (right panel) at each iteration for binaries with equal masses, vanishing spins, $a_{T}=15$ M (top) or $r_{a,T}=60$ M (bottom), and different values of target eccentricity $e_{T}$ (color bar). We achieve parameter convergence for all the cases we attempted, including eccentricities up to $0.65$. For $e_{T}>0.65$, the BHs risk head-on collision, which terminates the control loop prematurely before obtaining $\vec{\theta}_{ID,T}$. Thus we omit cases $e_{T}>0.65$ in this work. As $e_{T}$ increases to 0.65, more iterations are required for parameter convergence. We also inlay separation over time plots based on the final (most converged) iteration of the control loop.
The remaining panels show the difference between the achieved (fitted from the NR data) and the target eccentricity (top right), semimajor axis (bottom left), and mean anomaly (bottom right) as a function of iteration. Here and in subsequent figures, $a^{(i)},e^{(i)}$ and $\ell^{(i)}$ refer to the fitted orbital parameters in each iteration $i$ of the control loop. Even though we only impose a tolerance threshold on the eccentricity, all orbital parameters converge to their target values after $3$ iterations. In fact, the mean anomaly, $\ell$, converges almost immediately. Since this behavior is typical in subsequent explorations, we omit the mean anomaly from now on. Moreover, the semimajor axis converges in 1 fewer iteration than the eccentricity. This justifies continuing to consider only the eccentricity in the parameter control tolerance.
@@ III.1 Impact of target eccentricity value
Figure 5 shows parameter convergence at fixed $a_{T}$ and $e_{T}<0.5$ (top) and fixed $r_{a,T}$ and $e_{T}\geq 0.5$ (bottom). We show the difference between the target and fitted eccentricity (left panels) and semimajor axis (right panels) as a function of iteration and colored by the target eccentricity value. The inlaid plots show the BH coordinate separation over time, using the same color scheme. In the top row, the separation reflects a constant $a_{T}$ for each $e_{T}<0.5$. As the eccentricity increases (darker purple), the average separation stays constant but the apastron and periastron passages become more extreme. In the bottom row, all trajectories begin at the same apastron and result in roughly similar periastron values.
In all cases, i.e., for target eccentricty up to $e_{T}=0.65$, the eccentricity and semimajor axis converge to their target value, with higher eccentricities requiring on average more iterations. This is likely because higher eccentricity orbits are more relativistic, thus the post-Newtonian approximation to the dynamics is less accurate. Such less accurate fitting equations lead to less efficient initial data parameter updates in each control loop iteration, thus requiring more iterations to converge.
@@ III.2 Aligned spins
Figure 6: Impact of aligned spins on parameter control. We plot the difference between the fitted and target eccentricity (left panel) and semimajor axis (right panel) at each iteration for binaries with equal masses, $a_{T}=15$ M, $e_{T}=\{0.1,0.2,0.3\}$ (top to bottom) and different values of the aligned spin (color bar). We successfully converge to the target parameters for eccentric, aligned spin binaries, with higher eccentricities again requiring more iterations.
Figure 7: Comparing the simultaneous (top) and the sequential (bottom) parameter control approaches. We consider equal-mass binaries with $a_{T}=15$ M, $e_{T}=0.1$, and different values of $\chi_{1,x,T}$ (color bar). We show the difference between the target and the achieved value for the eccentricity (left), semimajor axis (middle), and in-plane spin angle, $\arccos(\vec{\chi}_{T}\cdot\vec{\chi}^{(i)})$ (right). In the sequential case, we plot results from the second parameter control loop, i.e., after obtaining a simulation with the target orbital parameters and no in-plane spins. Both methods achieve convergence, though as expected the simultaneous method needs more iterations than the second parameter control loop for the simultaneous method. Moreover, the orbital parameters are closer to their target value in the first iteration for the sequential method (bottom) and the spin direction converges monotonically (right).
Next, we explore the impact of aligned spins. Aligned spins remain constant in both direction and magnitude (modulo horizon absorption effects Scheel et al. (2015)). Therefore the rotations of Fig. 3 reduce to the identity, and the spins are not updated during the eccentricity-control iterations. However, the spins still affect the binary dynamics and rate of orbital decay and could therefore impact the orbital parameter convergence. We study binaries with equal masses, $a_{T}=15$ M, $e_{T}=\{0.1,0.2,0.3\}$, and vary the amount of aligned spin $\chi_{z,A,T}=\chi_{z,B,T}$ in $[-0.7,0.7]$. Results are shown in Fig. 6 for the different target eccentricities (top to bottom) and aligned spin values (color bar). In all cases, the eccentricity and semimajor axis converge to their target value after $\leq 7$ iterations.
@@ IV Eccentric, precessing orbits
In this section, we generalize to precessing binaries. Going forward, we assume one spinning BH and no aligned component for simplicity. In-plane spins precess in the inertial frame, causing the BH frame to change orientation. The parameter control loop of Fig. 2 and the updating formulas in Eqs. (11)–(13) and (14) treat the orbital parameters and spins independently for convenience. Since the orbital parameters and the spins are updated independently, we begin by exploring two parameter control methods:
@@
    Simultaneous. We follow the process of Fig. 2 where in the first iteration we simulate an eccentric, precessing binary and then iteratively update the orbital parameters and the spins, per Eqs. (11)–(13) and (14).
@@
Figure 7 compares parameter convergence for the simultaneous (top) and the sequential (bottom) methods for an equal-mass binary, $a_{T}=15$ M, $e_{T}=0.1$ and different values of $\chi_{1,x,T}$, and with $\chi_{1,y,T}=\chi_{1,z,T}=|\vec{\chi}_{2,T}|=0$. For the sequential case, we present only the second parameter control loop, i.e., the one that updates both the orbital parameters and the spins. Beyond the eccentricity (left) and the semimajor axis (middle), we show the angle between the target and the actual spin at $t_{T}$ (right). As before, all parameter control loops use a tolerance threshold based solely on eccentricity.
@@
Figure 8: Parameter convergence for precessing spin and varying spin magnitudes. We plot the difference between the target and the fitted eccentricity (left), semimajor axis (middle), and spin direction (right) for different values of the magnitude and direction of the in-plane spin. We choose $|\vec{\chi}_{1,T}|=0.3$ (solid) and $|\vec{\chi}_{1,T}|=0.6$ (dashed). The angle between the target in-plane spin and the $x$-axis is denoted as $\arccos(\vec{\chi}_{x}\cdot\vec{\chi}_{T})$ (color bar). In all cases, the parameters converge after $\leq 6$ iterations with behavior depending more on the in-plane spin magnitude than the in-plane spin angle, c.f. also the top panel of Fig. 7, which is similar to this plot except it compares different in-plane spin magnitudes with the same in-plane spin angle.
Comparing the rate of parameter convergence between the two methods, the sequential method results in orbital parameters closer to their target values at the first iteration. This is evident for both the eccentricity and semimajor axis. Even though the second control loop in the sequential method plotted here starts from preconverged orbital parameters, they still differ from their target value. This is because the first loop achieved the target parameters under zero in-plane spins, which are added in the second loop. Therefore the difference between the actual and the target eccentricity and semimajor axis in the first iteration is a consequence of the impact of in-plane spins on orbital dynamics.
Simultaneously adjusting the orbital parameters and the spin directions (top) results in spins that are no longer monotonically converging to their target value. Between the first and the second iteration, and even though both the eccentricity and the semimajor axis are closer to their target values, the spin angle difference increases from $\sim 5-10\,$deg to $\sim 20-30\,$deg. This is again a consequence of the coupling of orbital and spin dynamics. The spin is updated per Eq. (14) for the second iteration assuming the orbit of the first iteration. Despite this, in the third iteration where the orbit is not substantially updated, the spin angle difference quickly decreases. Overall, we achieve differences between the target and the actual spin angle ${\cal{O}}(\rm{deg})$, even though spins are not considered in the tolerance threshold.